Find the inverse of
f(x)= 4-x divded by x+6
a. Sate the domain
b. State the domain of f1(x)
c. State the range of f(x)
: Find the inverse of
f(x)= 4-x divded by x+6
a. Sate the domain
b. State the domain of f1(x)
c. State the range of f(x)

Answer by jasmine4195(3)

I think that means, that if you have:
y = (6x + 4)/(4x + 5)
You want to solve for x in terms of y. so:
(4x+5)y = 6x+4
4xy + 5y = 6x + 4
4xy - 6x = 4 -5y
x(4y - 6) = 4 - 5y
x = (4 - 5y) / (4y - 6) and if you have to like simplify it this is a good way to do it
Jasmine4195